• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi/chiranjib/Jin

    Log entry time 22:24:39 on February 05, 2009

    Entry number 259124

    keyword=Laser alignment status

    Laser alignment is almost done.
    Fiber 04,05 and N6 are connected to the longitudinal line. Now the lasers
    are in full power. We plan to try out EPR now.